Q1.Consider an economy that produces only chocolate bars. In year 1, the quantity produced is 3 bars and the price is Ghc4. In year 2, the quantity produced is 4 bars and the price is Ghc5. In year 3, the quantity produced is 5 bars and the price is Ghc6. Year 1 is the base year.

(i) What is the nominal GDP for each of these years

(ii) What is the real GDP for each of these years?

iii) Calculate the GDP deflator for each period

(iv) What is the % growth rate of real GDP from year 1 to year 2?

(v) What is the % growth rate of nominal GDP from year 1 to year 3



1
Expert's answer
2021-07-01T13:49:43-0400

(i)

"nominal \space GDP \\in\space year 1= 3\times 4=Ghc\space 12\\ in\space year 2= 4\times 5= Ghc\space 20\\ in \space year 3= 5\times 6= Ghc\space 30"


(ii)

"Real \space GDP \\ year 1= \$3\\ year 2= \$4\\ \\ year 3=\$5"


iii)

"GDP \space Deflator\\ in\space year 1= \frac{12}{3}= 4.\\ year 2=\frac{ 20}{4} =5 \\ year 3=\frac{ 30}{5}= 6"


(iv)

"=\frac{4-3}{3}=33.33\%"


(v)

"=\frac{30-12}{12}=175\%"